OpenStudy (anonymous):

A circle has a radius of 11 cm. What is the approximate area of the circle? Use 3.14 to approximate pi. Write the answer to the nearest hundredth.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

multiply 3.14*11^2

OpenStudy (anonymous):

then round that to the nearest hundredth

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

multiply 3.14 by 11 to the second power to get your answer

OpenStudy (igreen):

\(A = \pi r^2\) \(A = (3.14)(11^2)\) Simplify

OpenStudy (anonymous):

379.94

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Correct.

OpenStudy (anonymous):

so 379.94 is my answer

OpenStudy (igreen):

Yep